OpenStudy (anonymous):

How many thousands of gallons of water are there in a pond that has a circular diameter of 65’ and an average depth of 145” ? (Round up to the nearest thousand of gallon)

OpenStudy (anonymous):

Start by finding the volume: V=pi*r^2*h = pi*(65/2 ft)^2*145/12 ft = 1233.7 ft^3 Then convert from ft^3 to gallons: Google says

OpenStudy (anonymous):

so 9 thousand gallons seems like its your answer